write the point slope form of the equation of the line that passes through the point and has the given slope, then rewrite the equation in slope-intercept form
(-2, 4), m = 3
Thanks for help

Answer by checkley71(8403)

Y=mX+b WHERE m=SLOPE & b=Y INTERCEPT.
NOW SUBSTITUTE THE X&Y VALUES & THE SLOPE (m) & SOLVER FOR b.
4=3*-2+b
4=-6+B
b=4+6
b=10 THUS WE HAVE THE FOLLOWING EQUATION.
Y=3X+10